A 0.43-kg metal sphere oscillates at the end of a vertical spring. As the spring stretches from 0.12 m to 0.23 m (relative to its unstrained length), the speed of the sphere decreases from 7.2 to 4.8 m/s. What is the spring constant of the spring?
Answer #1

Take gravitational potential energy when the spring is stretched by 0.23 m as 0
When the spring is stretched by 0.12 m, then it is at height h = 0.23 - 0.12 = 0.11 m above the point when it is stretched by 0.23 m

By conservation of energy
m*g*h1 + 0.5*k*x1^2 + 0.5*m*v1^2 = 0 + 0.5*k*x2^2 + 0.5*m*v2^2
0.43*9.8*0.11 + 0.5*k*0.12^2 + 0.5*0.43*7.2^2 = 0.5*k*0.23^2 + 0.5*0.43*4.8^2
k = 345.742 N/m